What is the simplified form of cot70+4cos70
A) 1/root3 b)root3. C) 2root3. D) 1/2

Grade:11

1 Answers

Arun
5 years ago
cot70 + 4cos70= (cos 70 )/(sin 70) + 4cos70= cos 70 (1/sin 70) + 4We cannot go any further without the help of tables. Taking the help of tables for the sine values, the above expression reduces to0.34 (1/0.94 + 4)=1.73 (approx)= √3